$8,600 is invested for 5 months at 9% per year what is the future value
Answer by mananth(16949)

I= p*r*t
p= 8600
r=9% ( 0.09)
t=5/12 years
..
I= 8600 *0.09 *5/12
I=322.5
Future value = $8600+322.5 = $8922.50